OpenStudy (anonymous):

You need a piece of fabric 5 12 yard long. Which of these is the smallest piece of fabric that would work?

OpenStudy (anonymous):

Now I know the answer is 1/2 a yard... Can someone please explain why?

OpenStudy (anonymous):

Is this a fraction? 5/12 ? I assume there is a list of choices? 6/12 would be 1/2.

OpenStudy (anonymous):

Yes. 5/12. The choices are 1/2, 2/3 and 5/16 yards.

OpenStudy (anonymous):

The only way I would evaluate it is the fraction 5/12 = 0.41. 1/2 = 0.5, 2/3 = 0.66, and 5/16 = 0.31. So the only choice that would work according to the criteria is 1/2.

OpenStudy (anonymous):

Because it's the closest value to 5/12?